The equation of line r is y-2=1/10(x+9). line "s" is parallel to line "r" and passes through (10,1) What is the equation of line "s"?

Write the equation in slope-intercept form. Write the numbers in the equation as proper fractions, improper fractions, or integers.

Answer by mananth(16946)

The equation of line r is y-2=1/10(x+9). line "s" is parallel to line "r" and passes through (10,1) What is the equation of line "s"?

line r
y-2=1/10(x+9)
y= 1/10(x+9) +2
y=(1/10)x +9/10 +2
y= (1/10)x+29/2
slope of line = 1/10
line s is parallel to line r
slope will be same = 1/10
point (10,1)
equation will be
(y-1)=1/10(x-10)
y-1=(1/10)x-1
y= (1/10)x